The rate law for the reaction NH4+(aq) + NO2−(aq) → N2(g) + 2H2O(l) is given by rate = k[NH4+][NO2−]. At a certain temperature, the rate constant is 3.80 × 10−4 /M·s. Calculate the rate of the reaction at that temperature if [NH4+] = 0.261 M and [NO2−] = 0.160 M. × 10 M/s Enter your answer in scientific notation.
Answer #1

Rate=k*[NH4+]*[NO2-]

[NH4+] = 0.261 M ,[NO2−] = 0.160 M

rate constant=k= 3.80 × 10-4/M·s.

Rate=3.80*10-4*0.261*0.160=1.587*10-5M/s.